An 8 kg package in a mail-sorting room slides 2 m down a chute that is inclined at 53 degrees below the horizontal. The coefficient of kinetic friction between the package and the chute's surface is 0.4.

Calculate the work done on the package by...
A) Friction
B) Gravity
C) The normal force
D) What is the net work done on the package?
Answer #1
Concepts and reason

The main concept used to solve the problem is relationship between work done, force and distance.

Initially, determine the forces acting on the block. Later, find the displacement of the block along the direction of force. Finally, multiply the force and displacement to calculate the work done.

Fundamentals

The work done on an object is,

W=FdW = F \cdot d

Here, WW is the work done, FF is the force, and dd distance.

(A)

The work done due to friction is,

Wfriction=Ffrictionald{W_{friction}} = {F_{frictional}} \cdot d

Here, Wfriction{W_{friction}} is the work done due to frictional force, Ffrictional{F_{frictional}} is the frictional force, and dd is the displacement due to the frictional force.

The frictional force is,

Ffrictional=μkN{F_{frictional}} = - {\mu _k}N

Here, μk{\mu _k} is the coefficient of friction and NN is the normal force.

The negative sign indicates that the frictional force acts in the opposite direction of the block’s motion.

The normal force is,

N=mgcosθN = mg\cos \theta

Here, mm is the mass and ggis the acceleration due to gravity, and θ\theta is the inclination.

Substitute N=mgcosθN = mg\cos \theta in Ffrictional=μkN{F_{frictional}} = {\mu _k}N to calculate the frictional force. The frictional force is,

Ffrictional=μk(mgcosθ)=μkmgcosθ\begin{array}{c}\\{F_{frictional}} = - {\mu _k}\left( {mg\cos \theta } \right)\\\\ = - {\mu _k}mg\cos \theta \\\end{array}

Substitute Ffrictional=μkmgcosθ{F_{frictional}} = - {\mu _k}mg\cos \theta in Wfriction=Ffrictionald{W_{friction}} = {F_{frictional}} \cdot d to calculate the work done due to frictional force.

The work done is,

Wfriction=(μkmgcosθ)d=μkmgdcosθ\begin{array}{c}\\{W_{friction}} = \left( { - {\mu _k}mg\cos \theta } \right) \cdot d\\\\ = - {\mu _k}mgd\cos \theta \\\end{array}

Substitute 0.40.4 for μk{\mu _k}, 8.0kg8.0\,{\rm{kg}} for mm, 2.0m2.0\,{\rm{m}} for dd, 9.8m/s29.8\,{\rm{m/}}{{\rm{s}}^2} for gg, and 5353^\circ for θ\theta .

The work done is,

Wfriction=(0.4)(8.0kg)(9.8m/s2)(2.0m)(cos(53))=37.7J\begin{array}{c}\\{W_{friction}} = - \left( {0.4} \right)\left( {8.0\,{\rm{kg}}} \right)\left( {9.8\,{\rm{m/}}{{\rm{s}}^2}} \right)\left( {2.0\,{\rm{m}}} \right)\left( {\cos \left( {53^\circ } \right)} \right)\\\\ = - 37.7\,{\rm{J}}\\\end{array}

The work done due to friction is 37.7J - 37.7\,{\rm{J}}.

The negative sign indicates that the energy is lost due to work done by friction.

(B)

The work done due to gravity is,

Wgravity=Fgravitationald{W_{gravity}} = {F_{gravitational}} \cdot d

Here, Wgravity{W_{gravity}} is the work done due to gravitational force, Fgravitational{F_{gravitational}} is the gravitational force and dd is the displacement due to the gravitational force.

The gravitational force is,

Fgravitational=mgsinθ{F_{gravitational}} = mg\sin \theta

Here, mm is the mass, θ\theta is the inclination and gg is the acceleration due to gravity.

Substitute Fgravitational=mgsinθ{F_{gravitational}} = mg\sin \theta in Wgravitational=Fgravitationald{W_{gravitational}} = {F_{gravitational}} \cdot d to calculate the work done due to gravitational force. The work done is,

Wgravitational=(mgsinθ)d=mgdsinθ\begin{array}{c}\\{W_{gravitational}} = \left( {mg\sin \theta } \right) \cdot d\\\\ = mgd\sin \theta \\\end{array}

Substitute 8.0kg8.0\,{\rm{kg}} for mm, 2.0m2.0\,{\rm{m}} for , 5353^\circ for θ\theta , and 9.8m/s29.8\,{\rm{m/}}{{\rm{s}}^2} for gg.

Wgravitational=(8.0kg)(9.8m/s2)(2.0m)(sin(53))=125.2J\begin{array}{c}\\{W_{gravitational}} = \left( {8.0\,{\rm{kg}}} \right)\left( {9.8\,{\rm{m/}}{{\rm{s}}^2}} \right)\left( {2.0\,{\rm{m}}} \right)\left( {\sin \left( {53^\circ } \right)} \right)\\\\ = `125.2\,{\rm{J}}\\\end{array}

The work done due to gravity is 125.2J125.2\,{\rm{J}}.

(C)

The work done due to the normal force is,

Wnormal=Nd{W_{normal}} = N \cdot d

Here, Wnormal{W_{normal}} is the work done due to normal force, NN is the normal force and dd is the displacement due to the normal force.

The displacement due to normal force is zero. Substitute 00 for dd. The work done is,

Wnormal=N(0)=0N\begin{array}{c}\\{W_{normal}} = N\left( 0 \right)\\\\ = 0\,{\rm{N}}\\\end{array}

The work done due to normal force is 0J0\,{\rm{J}}

(D)

The net work done is,

W=Wfrictional+Wgravitational+WnormalW = {W_{frictional}} + {W_{gravitational}} + {W_{normal}}

Substitute 37.7J - 37.7\,{\rm{J}} for Wfrictional{W_{frictional}}, 0J0\,{\rm{J}} for Wnormal{W_{normal}} and 125.2J125.2\,{\rm{J}} for Wgravitational{W_{gravitational}}. The net work done is,

W=37.7J+0J+125.2=87.5J\begin{array}{c}\\W = - 37.7\,{\rm{J}} + 0\,{\rm{J}} + 125.2\\\\ = 87.5\,{\rm{J}}\\\end{array}

The net work done on the block is 87.5J87.5\,{\rm{J}}.

Ans: Part A

The work done due to frictional force is 37.7J - 37.7\,{\rm{J}}.

Part B

The work done due to gravity is 1.2×102J1.2 \times {10^2}\,{\rm{J}}.

Part C

The work done due to normal force is 0J0\,{\rm{J}}

Part D

The net work done on the block is 87.5J87.5\,{\rm{J}}.